Find values for m and b in the following system so that the solution to the system
is (-3, 4).
5x+7y=b
mx+y=22

Answer by edjones(8007)

5x+7y=b
mx+y=22
.
5*-3 + 7*4
=-15+28
=13
b=13
.
-3m*+4=22
-3m+4-4=22-4
-3m=18
-3m/-3=18/-3
m=-6
.
Ed